# Expanding RL with Verifiable Rewards Across Diverse Domains

### v1

@@ Abstract
Reinforcement learning (RL) with verifiable rewards (RLVR) has shown promising results in mathematical reasoning and coding tasks where well-structured reference answers are available. However, its applicability to broader domains remains underexplored. In this work, we study the extension of RLVR to more diverse domains such as medicine, chemistry, psychology, and economics. We observe high agreement in binary judgments across different large language models (LLMs) when objective reference answers exist, which challenges the necessity of large-scale annotation for training domain-specific reward models. To address the limitations of binary rewards when handling unstructured reference answers, we further incorporate model-based soft scoring into RLVR to improve its flexibility. Our experiments show that a distilled generative reward model can serve as an effective cross-domain verifier, providing reliable reward signals for RL without requiring domain-specific annotations. By fine-tuning a base 7B model using various RL algorithms against our reward model, we obtain policies that outperform state-of-the-art open-source aligned LLMs such as Qwen2.5-72B-Instruct and DeepSeek-R1-Distill-Qwen-32B by a large margin, across domains in free-form answer settings. This also strengthens RLVR’s robustness and scalability, highlighting its potential for real-world applications with noisy or weak labels.
@@ 1 Instruction
Reinforcement learning with verifiable rewards (RLVR) has emerged as a promising training strategy for improving the reasoning capabilities of large language models (LLMs) (Luong et al., 2024; Lambert et al., 2024), even in the absence of a supervised fine-tuning stage (Guo et al., 2025). This is applied in a reference-based setting, where training assumes access to an objective ground-truth label for each prompt. A reward or verification function typically assigns a binary reward, indicating whether a model’s response aligns with the ground truth.
Previous research has primarily demonstrated the effectiveness of this method in tasks with well-structured answers in limited domains, such as mathematical reasoning and coding. In these settings, reference answers such as numerical values or predefined test cases enable straightforward rule-based verification. To improve the robustness of verifiers across tasks, recent studies have also explored training reward models within a single domain. For instance, researchers utilize approximately 800k instances to train a reward model for mathematical reasoning (Team et al., 2025).
In this work, we extend the study of RL with verifiable rewards beyond well-structured mathematics and coding tasks, investigating its potential to improve LLMs’ complex reasoning abilities across diverse domains, such as medicine, chemistry, psychology, economics, and education. Since training a separate reward model for each domain is impractical, we begin by querying an aligned, general-domain LLM, inspired by previous reference-free studies that demonstrated the superiority of generative verifiers (Zhang et al., 2024a; Mahan et al., 2024) over discriminative reward models in selecting the highest-scoring response from multiple samples during inference. Surprisingly, when provided with objective reference answers written by domain experts, we observe high agreement in binary judgments between closed-source and open-source LLMs (e.g., GPT-4o vs. Qwen2.5-72B-Instruct (Team, 2024)) across diverse domains. This high consistency suggests that, when high-quality, objective reference answers are available, reference-based grading is generally easier than reference-free assessment, which is inherently as difficult as identifying the first mistake in a response (Lightman et al., 2023). As a result, the necessity of large-scale annotation or synthesis of training data for single-domain reward models comes into question.
Furthermore, previous RLVR studies have predominantly relied on binary reward scores (Gandhi et al., 2024; Zhang et al., 2024b; Lambert et al., 2024; Guo et al., 2025; Ma et al., 2025a), raising concerns about their robustness in complex reasoning tasks across diverse domains. Even within mathematical problem-solving, our analysis of real-world examination problems shows that only $60.3\%$ of reference answers are single-word numerical values or expressions that can be relatively easily verified using rule-based reward functions. This proportion decreases further to $45.4\%$ for multi-domain problems. These observations underscore the limitations of binary scoring in assessing a broader range of reasoning-intensive tasks and highlight the potential benefits of exploring more expressive scoring mechanisms.
To address this, we incorporate model-based soft reward scores into RLVR by leveraging the probability of a single token that represents a generative verifier’s final judgment. Noticeably, we demonstrate that when objective reference answers are available, a reasonably effective verifier can be trained for diverse domains using a relatively small LLM, such as a 7B model, while achieving downstream performance comparable to much larger off-the-shelf generative verifiers, such as Qwen2.5-72B-Instruct (Team, 2024). In particular, this is achieved without the need for task-specific annotations for reward models. Instead, we collect policy’s responses and their assessments from a teacher verifier during online exploration. These judged responses, with varying performance levels and formatting noise, may further enhance the robustness of the learned reward model (Section 3.3). By fine-tuning a base 7B model using three RL algorithms guided by our reward model, we obtain policies that outperform powerful open-source aligned LLMs – such as Qwen2.5-72B-Instruct (Team, 2024) and DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025) – by up to $8.0\%$ in accuracy across various domains in free-form answer tasks. In addition to improving generalization, model-based rewards scale effectively with increasing data size. In contrast, rule-based rewards struggle, especially when addressing unstructured reference answers, and may even lead to performance degradation.
Our main contributions are as follows.
  • We extend RLVR to diverse domains, demonstrating its effectiveness beyond mathematics and coding, particularly in handling unstructured reference answers.
  • We show that even with unstructured objective answers, a relatively small cross-domain reward model can be trained without requiring task-specific annotations. [^1] Our reward model and free-form data will be released at https://huggingface.co/collections/virtuoussy/rlvr-67ea349b086e3511f86d1c1f.
  • We propose incorporating model-based soft rewards into RLVR and empirically demonstrate their superiority, scalability, and robustness over rule-based rewards across domains.
@@ 3 Method
Figure 1: Overview paradigm of RLVR with our cross-domain verifier.
@@ 4.7 Out-of-Distribution Evaluation
To further validate the effectiveness of our reward model, we conduct additional evaluations on two benchmarks: NaturalReasoning (Yuan et al., 2025) and WebInstruct (Yue et al., 2024). We compare the performance of the rule-based reward with our RM-7B using RLOO with binary reward. The base model is Qwen2.5-7B. The results are shown in Table 3. As can be seen from the table, the performance of RM-7B remains significantly superior to the rule-based reward on datasets from other domains. This demonstrates that our general-purpose reward model can extend to other domains while maintaining strong performance.
@@ 5 Discussions and Conclusions
In addition, we do not consider format-based rewards (Guo et al., 2025) in this work. We revisit the role of format-related constraints and rewards in this context. In prior work, pattern-based functions are often used for scoring, making it critical to guide LLMs to enclose their final answers in an easily parsed format. These extracted answers are then compared with the reference answers for verification and evaluation. In contrast, by reintroducing a reward model in RLVR without imposing any format constraints on reference answers or model responses, we reduce the need for extensive human effort in data standardization and pattern design.

### v2

# Crossing the Reward Bridge: Expanding RL with Verifiable Rewards Across Diverse Domains

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) has demonstrated significant success in enhancing mathematical reasoning and coding performance of large language models (LLMs), especially when structured reference answers are accessible for verification. However, its extension to broader, less structured domains remains unexplored. In this work, we investigate the effectiveness and scalability of RLVR across diverse real-world domains including medicine, chemistry, psychology, economics, and education, where structured reference answers are typically unavailable. We reveal that binary verification judgments on broad-domain tasks exhibit high consistency across various LLMs provided expert-written reference answers exist. Motivated by this finding, we utilize a generative scoring technique that yields soft, model-based reward signals to overcome limitations posed by binary verifications, especially in free-form, unstructured answer scenarios. We further demonstrate the feasibility of training cross-domain generative reward models using relatively small (7B) LLMs without the need for extensive domain-specific annotation. Through comprehensive experiments, our RLVR framework establishes clear performance gains, significantly outperforming state-of-the-art open-source aligned models such as Qwen2.5-72B and DeepSeek-R1-Distill-Qwen-32B across domains in free-form settings. Our approach notably enhances the robustness, flexibility, and scalability of RLVR, representing a substantial step towards practical reinforcement learning applications in complex, noisy-label scenarios.
Figure 1: Overview paradigm of RLVR with our cross-domain verifier.
@@ 1 Instruction
Reinforcement learning with verifiable rewards (RLVR) has recently emerged as an effective paradigm for improving the reasoning capabilities of large language models (LLMs) (Luong et al., 2024; Lambert et al., 2024), even in scenarios without supervised fine-tuning (Guo et al., 2025). RLVR typically leverages reference-based signals, assuming the availability of objective ground-truth answers to determine whether model responses align with reference outcomes. In prior studies, RLVR has mainly demonstrated success on tasks with precisely structured solutions, such as mathematical reasoning or code generation, where binary verification signals (correct or incorrect) can be reliably computed with simple rule-based verifiers (Team et al., 2025; Gandhi et al., 2024; Zhang et al., 2024b). Nonetheless, the extension of RLVR to broader, more nuanced domains remains largely unexplored, due primarily to the challenges associated with verifying complex, frequently unstructured reference answers.
In this paper, we aim to extend the applicability of RLVR to domains beyond structured mathematics and coding, by investigating its performance in a diverse set of complex reasoning-intensive areas such as medicine, chemistry, psychology, economics, and education. Central to this exploration is the observation that binary correctness judgments, even on broad-domain tasks, tend to exhibit remarkable agreement across varied large language models (LLMs), including both closed-source models (e.g., GPT-4o) and recently released powerful open-source solutions (e.g., Qwen2.5-72B-Instruct) when provided high-quality objective references authored by domain experts. This finding indicates that reference-based evaluation of diverse domain answers is typically easier than reference-free verification, which is inherently as difficult as identifying the first mistake in a response (Lightman et al., 2023). Consequently, this insight undermines the presumed necessity for extensive domain-specific annotation and motivates rethinking traditional practices in reward-model training for multi-domain scenarios.
While binary rewards have been the prevalent standard across RLVR applications (Gandhi et al., 2024; Lambert et al., 2024; Guo et al., 2025; Ma et al., 2025a), they pose clear limitations—especially for unstructured tasks. Notably, our data analysis on real-world exam questions reveals that only 60.3% of mathematical problems possess single-term numerical answers verifiable by rule-based methods, with the ratio dropping further to 45.4% for complex multi-domain queries. This presents inherent challenges for binary reward schemes and demonstrates the need for richer and more granular verification mechanisms. To address these limitations, we propose incorporating soft scores obtained from generative, model-based verifiers directly into RLVR. Specifically, we compute a soft reward from the probability of a single indicative token produced by a generative verifier summarizing its assessment. Crucially, we demonstrate that it is feasible to distill effective multi-domain generative verifier models based on relatively compact models (sizes as small as 7B) without conducting extensive domain-specific annotation. Instead, we employ data composed of response samples and their corresponding judgments collected during RL exploration under the supervision of a larger cross-domain generative teacher model. These noisy yet more realistic datasets promote robustness of the subsequently distilled model-based rewards.
Our empirical results strongly validate the effectiveness of our extended RLVR framework across various domains. By fine-tuning modest-sized (7B) base models using various RL algorithms and our soft reward verifier, we obtain improved reasoning policies superior to state-of-the-art open-source alignment models such as Qwen2.5-72B-Instruct and DeepSeek-R1-Distill-Qwen-32B, achieving performance boosts of up to 8.0% accuracy in diverse, free-form reasoning tasks. We particularly observe that our model-based soft rewards consistently scale better and produce more robust policies compared to conventional rule-based binary rewards, especially on unstructured answer scenarios and larger training data regimes.
Contributions. Our key contributions can be summarized as follows:
  • We extend reinforcement learning with verifiable rewards (RLVR) to diverse domains, establishing its effectiveness beyond traditional structured answer scenarios.
  • We introduce and validate a novel framework incorporating generative model-based soft rewards within RLVR, demonstrating substantial improvements in generalization, robustness, and scalability relative to traditional binary rule-based rewards.
  • We empirically demonstrate the feasibility and efficacy of training compact (7B-scale) cross-domain generative reward verifiers without extensive domain-specific annotation, challenging traditional assumptions about annotation scale.
  • We release a dataset containing 570k examples of multi-domain free-form data and the corresponding trained reward model, available at https://huggingface.co/collections/virtuoussy/rlvr-67ea349b086e3511f86d1c1f, to facilitate future research in this promising direction.
@@ 4.7 Out-of-Distribution Evaluation
To further validate the effectiveness of our reward model, we conduct additional evaluations on two benchmarks: NaturalReasoning (Yuan et al., 2025) and WebInstruct (Yue et al., 2024). We compare the performance of the rule-based reward with our RM-7B using RLOO with binary reward. The base model is Qwen2.5-7B. For both datasets, we randomly select 30K examples for training and 5K sample for evaluation. The results are shown in Table 3. As can be seen from the table, the performance of RM-7B remains significantly superior to the rule-based reward on datasets from other domains. This demonstrates that our general-purpose reward model can extend to other domains while maintaining strong performance.
@@ 5 Discussions and Conclusions
In addition, we do not consider format-based rewards (Guo et al., 2025; Xie et al., 2025) in this work. We revisit the role of format-related constraints and rewards in this context. In prior work, pattern-based functions are often used for scoring, making it critical to guide LLMs to enclose their final answers in an easily parsed format. These extracted answers are then compared with the reference answers for verification and evaluation. In contrast, by reintroducing a reward model in RLVR without imposing any format constraints on reference answers or model responses, we reduce the need for extensive human effort in data standardization and pattern design.